OpenStudy (anonymous):

A circle of radius r has a central angle of θ=15° which intercepts an arc of 27 inches. what is the value of radius r?

OpenStudy (anonymous):

2 *pi * r= circumference

OpenStudy (anonymous):

2 *pi is in radians form convert that to degrees you get 2*pi=360 degrees

OpenStudy (anonymous):

15°/360°= x x*2*pi * r= 27

OpenStudy (anonymous):

solve for x first then solve for r